OpenStudy (anonymous):

convert to an equation in rectangular coords: r=1/(2-cos(theta))

OpenStudy (anonymous):

cosθ = x/r r = 1/(2 - x/r)

OpenStudy (jdoe0001):

do you know what a "perfect trinomial square" is?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

But what does that have to do with this?

OpenStudy (jdoe0001):

hmm one sec

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ r^2=x^2+y^2\implies r=\sqrt{x^2+y^2}\qquad x=rcos(\theta)}}\\ --------------------------\\ r=\cfrac{1}{2-cos(\theta)}\implies r[2-cos(\theta)]=1\implies 2r-2cos(\theta)=1\\ \quad \\ 2(\sqrt{x^2+y^2})-x=1\implies 2(\sqrt{x^2+y^2})=x+1\\ \quad \\ \textit{squaring both sides}\\ \quad \\ 2^2(\sqrt{x^2+y^2})^2=(x+1)^2\implies 4(x^2+y^2)=x^2+2x+1\\ \quad \\ 4x^2+4y^2=x^2+2x+1\quad \textit{now grouping a few}\\ \quad \\ 3x^2+2x+4y^2=1\implies (3x^2+2x)+4y^2=1\)

OpenStudy (jdoe0001):

hmm I have a quick typo, dohh anyhow \(\bf{\color{blue}{ r^2=x^2+y^2\implies r=\sqrt{x^2+y^2}\qquad x=rcos(\theta)}}\\ --------------------------\\ r=\cfrac{1}{2-cos(\theta)}\implies r[2-cos(\theta)]=1\implies 2r-rcos(\theta)=1\\ \quad \\ 2(\sqrt{x^2+y^2})-x=1\implies 2(\sqrt{x^2+y^2})=x+1\\ \quad \\ \textit{squaring both sides}\\ \quad \\ 2^2(\sqrt{x^2+y^2})^2=(x+1)^2\implies 4(x^2+y^2)=x^2+2x+1\\ \quad \\ 4x^2+4y^2=x^2+2x+1\quad \textit{now grouping a few}\\ \quad \\ 3x^2+2x+4y^2=1\implies (3x^2+2x)+4y^2=1\)

OpenStudy (jdoe0001):

\(\bf (3x^2+2x)+4y^2=1\implies 3\left(x^2+\frac{2}{3}x+{\color{red}{ \square }}^2\right)+4(y-0)^2=1\) so... we would just need to make the 1st group, with the "x", a perfect square trinomial, thus we'd need a value that would give us the middle term

OpenStudy (jdoe0001):

if you notice the equation a bit, it'd look like an "ellipse" equation

OpenStudy (jdoe0001):

we know that the middle term will be , 2 times the square of the left and right terms so, what we have there is \(\bf \frac{2}{3}x\) that means that the other term will be \(\bf 2\cdot x\cdot {\color{red}{ \square }}=\cfrac{2}{3}x\implies {\color{red}{ \square }}= \cfrac{2\cancel{x}}{3\cdot 2\cdot \cancel{x}}\)

OpenStudy (jdoe0001):

so... what do you think is our trinomial?

OpenStudy (anonymous):

I'm still confused... why can't you stop at (3x2+2x)+4y2=1

OpenStudy (jdoe0001):

well, I guess you can but I assume you're expected to keep it in some "standard form", thus

OpenStudy (jdoe0001):

anyhow, if you don't need it as an ellipse standard form, then use -> (3x2+2x)+4y2=1 <- and that'd work

OpenStudy (anonymous):

Actually my prof did say something about completing the square. How did you get 2x/3⋅2⋅x?

OpenStudy (anonymous):

I'm very confused by this idea of polar coordinates.